The numerical values of the three sides of a triangle are given by three consecutive even integers. If the perimeter is greater than 30 inches, what are the possibilities for the shortest side?
The length of a rectangle is 4 times the width. If the perimeter is to be greater than or equal to 170 meters. What are the possible values for the width?
Please help me :(

Answer by ankor@dixie-net.com(22740)

The numerical values of the three sides of a triangle are given by three consecutive even integers. If the perimeter is greater than 30 inches, what are the possibilities for the shortest side?
:
let s = the shortest side
then
(s+2) = the next even integer
and
(x+4) = the one after that
:
x + (x+2) + (x+4) > 30
3x + 6 > 30
3x > 30 - 6
3x > 24
x > 24/3
x > 8
next even integer = 10 is one shortest side possibility
:
10 + 12 + 14 = 36
:
:
The length of a rectangle is 4 times the width.
If the perimeter is to be greater than or equal to 170 meters.
What are the possible values for the width?
Let w - the width
then
(w+4) = the length
2(w+4) + 2w >= 170
simplify, divide equation by 2
w + 4 + w >= 85
2w >= 85 - 4
2w >= 81
w >= 81/2
w >= 40.5 are the possible widths